# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection

### v2

@@ 6. Threats to Validity
Construct Validity: One primary threat to the construct validity of this study is the construction of the datasets. We first used DesigniteJava (designitejava, ) to detect potential positive and negative samples, then manually reviewed all the candidate data to validate the datasets, ensuring there were no false positives or negatives.
First, relying solely on DesigniteJava and PMD for initial sample detection may introduce bias. However, DesigniteJava and PMD are well-established and widely adopted tools in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ; tushar2016does, ; taibi2017how, ), providing a robust foundation for the identification of potential code smells. We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
@@
Lastly, due to budget constraints for commercial API usage, our evaluation of general-purpose LLMs (GPT-4o-mini and DeepSeek-v3) under ICL settings in RQ4 was limited to a maximum of three example pairs. This sample size may not fully represent the upper bounds of ICL performance. Future work could expand on this by increasing the number of example pairs and evaluating additional models to better understand the full potential of ICL performance. However, prior studies such as UniLog (xu2024unilog, ) have shown that the performance of LLMs on software engineering tasks does not always improve linearly with more examples, which suggests that while increasing the number of examples may be beneficial, performance gains are not guaranteed. As such, we believe our current evaluation still provides valuable insights and that this threat to validity is partially mitigated.
@@
Furthermore, we only employed four widely used and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. While these methods cover the major categories defined in the PEFT library, the exclusion of other approaches may limit the generalizability of our findings. For example, methods such as adapter tuning (houlsby2019parameter, ) could be explored in future work to provide a more comprehensive evaluation.
@@
Another potential threat to internal validity stems from conducting only a single experimental run per configuration due to computational resource constraints. Although we ensured consistency across all experiments by using a fixed random seed (42), this approach cannot fully account for the inherent stochasticity in the training process. To assess the impact of randomness, we conducted supplementary experiments using two alternative seeds (0 and 999) for a representative case (GraphCodeBERT on CC detection). The results shown in Table 11 indicate that there are only small variations in the performance metrics across different random seeds. These minor differences suggest that our core findings are robust to random initialization effects, with the overall trends remaining consistent. While some performance fluctuations are observed, they do not significantly alter the conclusions drawn from the experiments. However, we acknowledge that conducting multiple runs with different seeds would yield more statistically reliable conclusions.
@@
Additionally, data leakage represents another potential threat to internal validity. First, the poor zero-shot performance of GPT-3.5-turbo in detecting CC and CM code smells suggests that the model was not exposed to the test data (see Section 4.4), as we would expect stronger zero-shot effectiveness if GPT-3.5-turbo had memorized the data. Second, the significant improvements in performance for both small and large LMs following the application of PEFT techniques further suggest that memorization is unlikely to be an issue. Such substantial performance gains would not occur if the models had already memorized the test sets. Furthermore, Cao et al. (cao2024concerned, ) highlight that code LMs do not consistently exhibit reduced performance on data extending beyond their training cut-off dates. This careful management of data, along with the observed performance improvements after fine-tuning, ensures that data leakage was not a major threat, thereby maintaining the internal validity of our findings.
7. Related Work
7.1. Code Smell Detection
Existing approaches to code smell detection predominantly rely on heuristic-based methods, which utilize predefined rules to identify problematic code patterns (pmd, ; designitejava, ; campbell2013sonarqube, ; fokaefs2007jdeodorant, ; vidal2015jspirit, ; moha2010decor, ). These methods aim to flag potential smells by analyzing the source code against established standards. For instance, PMD (pmd, ), a widely used static code analysis tool, supports 16 programming languages and applies rule sets to detect code smells. These rule sets can be tailored to suit specific project requirements or design standards, making the tool adaptable for different contexts. Similarly, Sharma (designitejava, ) introduced DesigniteJava, an advanced tool capable of identifying code smells at multiple levels — ranging from architecture and design to implementation.
Traditional heuristic-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection. In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives. Maiga et al. (maiga2012support, ) implemented SVMDetect based on an ML technique — support vector machines. Fontana et al. (fontana2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
Liu et al. (liu2021deep, ) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not). Sharma et al. (sharma2021code, ) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells. Their work confirms the feasibility of applying transfer-learning for code smell detection. Ho et al. (ho2023fusion, ) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells. They conducted an empirical study to compare DeepSmells with other well-established detection tools. Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance. They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection.
Recently, researchers have started to explore the use of LLMs for code smell detection. Mesbah et al. (mesbah2025leveraging, ) used the MLCQ dataset to analyze how GPT-4 and LLaMA perform when prompted to identify and classify code smells. Silva et al. (silva2024detecting, ) employed two types of prompts to query ChatGPT for identifying Blob, Data Class, Feature Envy, and Long Method code smells. Sadik et al. (sadik2025benchmarking, ) created a dataset for evaluating LLMs on automated code smell detection. They benchmarked two state-of-the-art LLMs and found that LLMs are not yet ready to replace deterministic tools in all aspects of code quality assurance. Wu et al. (wu2024ismell, ) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection. They subsequently utilized the detection results to prompt LLMs for refactoring.
7.2. PEFT Methods on Software Engineering Tasks
Several studies have evaluated PEFT methods on various Software Engineering (SE) tasks. Wang et al. (wang2023one, ) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization. According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting. Liu et al. (liu2023empirical, ) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards, ) on both the code understanding task and code generation task. Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements. Liu et al. (liu2024delving, ) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation. Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension. Weyssow et al. (weyssow2023exploring, ) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation. Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse SE scenarios. Li et al. (li2024comprehensive, ) created an instruction dataset of Automated Program Repair (APR) using prompt engineering. They employed this dataset to fine-tune four LLMs using four different PEFT methods. The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
8. Conclusions and Future Work
In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection. We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data. Our analysis reveals that PEFT methods are more efficient than full fine-tuning and In-Context Learning (ICL) in most cases, underscoring the potential of PEFT methods to enhance code smell detection and offer valuable insights for optimizing LMs for other Software Engineering (SE) tasks. Additionally, we have constructed high-quality datasets for four widely studied code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
Data Availability
Acknowledgements.

### v3

@@ 7. Threats to Validity
Construct Validity: One primary threat to the construct validity of this study is the construction of the datasets. We first used DesigniteJava (designitejava) to detect potential positive and negative samples, then manually reviewed all the candidate data to validate the datasets, ensuring there were no false positives or negatives.
First, relying solely on DesigniteJava and PMD for initial sample detection may introduce bias. However, DesigniteJava and PMD are well-established and widely adopted tools in Java code smell detection (chaniotaki2021architecture; sharma2021qscored; shah2023mining; tushar2016does; taibi2017how), providing a robust foundation for the identification of potential code smells. We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
@@
Lastly, due to budget constraints for commercial API usage, our evaluation of general-purpose LLMs (GPT-4o-mini and DeepSeek-v3) under ICL settings in RQ4 was limited to a maximum of three example pairs. This sample size may not fully represent the upper bounds of ICL performance. Future work could expand on this by increasing the number of example pairs and evaluating additional models to better understand the full potential of ICL performance. However, prior studies such as UniLog (xu2024unilog) have shown that the performance of LLMs on software engineering tasks does not always improve linearly with more examples, which suggests that while increasing the number of examples may be beneficial, performance gains are not guaranteed. As such, we believe our current evaluation still provides valuable insights and that this threat to validity is partially mitigated.
@@
Furthermore, our study focused exclusively on four widely used and representative PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) implemented within the Hugging Face PEFT library. Techniques such as adapter tuning (houlsby2019parameter) were excluded because they are not yet well supported by the library and would require additional external implementations, which could compromise the consistency of our experimental setup. As a result, our findings may not fully generalize to PEFT methods beyond the four examined ones. To mitigate this limitation, we carefully selected diverse PEFT methods covering the major families (prompt-based, LoRA-based, and $(IA)^{3}$) to ensure representativeness across categories. In future work, we plan to incorporate adapter tuning and other emerging PEFT techniques as they become more maturely integrated into the Hugging Face ecosystem, thereby enabling a broader and more comprehensive comparison.
@@
In our experiments, we used a batch size of 1 for all models due to computational resource constraints, which ensured consistency but could potentially affect convergence and performance. To evaluate this, we conducted supplementary experiments with different batch sizes, reported in Appendix A.1, which show that batch size has a limited impact on overall detection performance across PEFT methods, suggesting the main conclusions remain robust. Similarly, each configuration was run only once using a fixed random seed (42), which does not fully capture the stochasticity inherent in training. Additional runs with two alternative seeds (0 and 999) for a representative case (GraphCodeBERT on CC detection), reported in Appendix A.2, indicate minor performance variations, confirming robustness to random initialization. Nevertheless, we acknowledge that more comprehensive experiments with larger batch sizes and multiple seeds across more models could provide further insights and increase statistical reliability.
CodeBERT, GraphCodeBERT, and CodeT5 have a maximum input length of 512 tokens, requiring longer code samples to be truncated. This could potentially omit semantically important content relevant for code smell detection. We analyzed token length distributions and manually reviewed a subset of long samples (50 positives and 50 negatives per smell type) in Appendix B. The review showed that essential indicators of each code smell typically occur within the first 512 tokens, and truncated segments rarely contain critical information. Thus, while context window limitations represent a potential threat, their impact on overall evaluation results is likely minimal.
Additionally, data leakage represents another potential threat to internal validity. First, the poor zero-shot performance of GPT-4o-mini in detecting CC and CM code smells suggests that the model was not exposed to the test data (see Section 5.4), as we would expect stronger zero-shot effectiveness if GPT-4o-mini had memorized the data. Second, the significant improvements in performance for both small and large LMs following the application of PEFT techniques further suggest that memorization is unlikely to be an issue. Such substantial performance gains would not occur if the models had already memorized the test sets. Furthermore, Cao et al. (cao2024concerned) highlight that code LMs do not consistently exhibit reduced performance on data extending beyond their training cut-off dates. This careful management of data, along with the observed performance improvements after fine-tuning, ensures that data leakage was not a major threat, thereby maintaining the internal validity of our findings.
8. Related Work
8.1. Code Smell Detection
Existing approaches to code smell detection predominantly rely on heuristics-based methods, which utilize predefined rules to identify problematic code patterns (pmd; designitejava; campbell2013sonarqube; fokaefs2007jdeodorant; vidal2015jspirit; moha2010decor). These methods aim to flag potential smells by analyzing the source code against established standards. For instance, PMD (pmd), a widely used static code analysis tool, supports 16 programming languages and applies rule sets to detect code smells. These rule sets can be tailored to suit specific project requirements or design standards, making the tool adaptable for different contexts. Similarly, Sharma (designitejava) introduced DesigniteJava, an advanced tool capable of identifying code smells at multiple levels — ranging from architecture and design to implementation.
Traditional heuristics-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection. In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives. Maiga et al. (maiga2012support) implemented SVMDetect based on an ML technique — support vector machines. Fontana et al. (fontana2016comparing) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
Liu et al. (liu2021deep) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not). Sharma et al. (sharma2021code) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells. Their work confirms the feasibility of applying transfer-learning for code smell detection. Ho et al. (ho2023fusion) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells. They conducted an empirical study to compare DeepSmells with other well-established detection tools. Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance. They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection.
Recently, researchers have started to explore the use of LLMs for code smell detection. Mesbah et al. (mesbah2025leveraging) used the MLCQ dataset to analyze how GPT-4 and LLaMA perform when prompted to identify and classify code smells. Silva et al. (silva2024detecting) employed two types of prompts to query ChatGPT for identifying Blob, Data Class, Feature Envy, and Long Method code smells. Sadik et al. (sadik2025benchmarking) created a dataset for evaluating LLMs on automated code smell detection. They benchmarked two state-of-the-art LLMs and found that LLMs are not yet ready to replace deterministic tools in all aspects of code quality assurance. Wu et al. (wu2024ismell) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection. They subsequently utilized the detection results to prompt LLMs for refactoring.
8.2. PEFT Methods on Software Engineering Tasks
Several studies have evaluated PEFT methods on various Software Engineering (SE) tasks. Wang et al. (wang2023one) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization. According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting. Liu et al. (liu2023empirical) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards) on both the code understanding task and code generation task. Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements. Liu et al. (liu2024delving) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation. Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension. Weyssow et al. (weyssow2023exploring) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation. Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse SE scenarios. Li et al. (li2024comprehensive) created an instruction dataset of Automated Program Repair (APR) using prompt engineering. They employed this dataset to fine-tune four LLMs using four different PEFT methods. The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
9. Conclusions and Future Work
In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection. We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data. Our analysis reveals that PEFT methods are more efficient than full fine-tuning and In-Context Learning (ICL) in most cases, underscoring the potential of PEFT methods to enhance code smell detection and offer valuable insights for optimizing LMs for other Software Engineering (SE) tasks. Additionally, we have constructed high-quality datasets for four widely studied code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class, and we provided the replication package of this work online (replicationPackage) to facilitate future research and ensure the reproducibility of our study results.
Data Availability
Acknowledgements.
Appendix
Appendix A The Impact of Other Hyper-Parameters on PEFT Methods
In addition to the primary experimental settings discussed in the main text, we investigate the effects of two additional hyper-parameters on PEFT methods: batch size and random seed selection. Understanding the sensitivity of PEFT methods to these factors is crucial for assessing the stability and reproducibility of our results.
Specifically, we examine how varying batch sizes during training influences model efficiency and prediction accuracy. We also evaluate performance variations across multiple random seeds to ensure that our findings are not driven by a particular initialization.
A.1. Batch Size
Table 11. Results of GraphCodeBERT with PEFT methods and full fine-tuning on detecting CC across different batch sizes
Batch Size Method Precision Recall F1 MCC
1 Full FT 74.06% 74.02% 74.01% 48.08%
Prompt Tuning 65.49% 65.33% 65.24% 30.81%
Prefix Tuning 70.08% 69.89% 69.82% 39.97%
LoRA 73.15% 73.15% 73.15% 46.30%
$(IA)^{3}$ 72.50% 72.50% 72.50% 45.00%
8 Full FT 72.10% 72.07% 72.06% 44.16%
Prompt Tuning 63.20% 63.15% 63.12% 26.36%
Prefix Tuning 71.24% 70.98% 70.89% 42.22%
LoRA 73.76% 73.59% 73.54% 47.34%
$(IA)^{3}$ 72.01% 71.85% 71.80% 43.85%
16 Full FT 73.19% 72.83% 72.72% 46.02%
Prompt Tuning 63.48% 63.48% 63.47% 26.96%
Prefix Tuning 69.66% 69.02% 68.77% 38.68%
LoRA 73.59% 73.59% 73.59% 47.18%
$(IA)^{3}$ 71.65% 71.30% 71.19% 42.95%
32 Full FT 73.08% 72.61% 72.47% 45.69%
Prompt Tuning 63.27% 63.26% 63.25% 26.53%
Prefix Tuning 69.32% 68.48% 68.13% 37.79%
LoRA 73.46% 73.15% 73.06% 46.61%
$(IA)^{3}$ 70.66% 69.35% 68.85% 39.99%
Table 11 presents the performance of GraphCodeBERT with different PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$), as well as full fine-tuning, for detecting the Complex Conditional (CC) code smell across various batch sizes (1, 8, 16, and 32).
From Table 11, we observe that batch size has a limited impact on overall detection performance. Full fine-tuning and LoRA consistently achieve the highest MCC values across all batch sizes, while $(IA)^{3}$ also maintains strong performance. Prompt tuning generally yields lower scores, and prefix tuning falls in between. Minor fluctuations are present; for example, LoRA attains slightly higher F1 and MCC scores at batch size 8 compared to batch size 1. Overall, the performance trends remain stable, suggesting that PEFT methods are relatively robust to batch size variations and that the main experimental conclusions are unlikely to be materially affected.
@@ A.1. Batch Size
A.2. Random Seed
Table 12 presents the results of GraphCodeBERT with different PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$), as well as full fine-tuning, for detecting the Complex Conditional (CC) code smell across three different random seeds (0, 42, and 999).
As shown in Table 12, the variations in performance metrics across different random seeds are minor. These small differences indicate that our core findings are robust to the effects of random initialization, with overall trends remaining consistent. While some fluctuations in performance are observed, they do not materially alter the conclusions drawn from the experiments. Nevertheless, we acknowledge that conducting multiple runs with different seeds would yield more statistically reliable results.
Appendix B The Impact of Context Window Limit to SLMs
In the process of data construction, considering our GPU memory limitations, we filtered out any samples that exceed 1,024 tokens. However, several SLMs used in our study, including CodeBERT, GraphCodeBERT, and CodeT5, can only accept input sequences up to 512 tokens (see Section 4.3). Consequently, any code sample longer than 512 tokens must be truncated before being fed into these models, which may lead to the loss of semantically important information and potentially affect the accuracy of code smell detection.
To quantify this effect, Figure 10 presents the token length distributions for CodeBERT and CodeT5 across different code smell categories. GraphCodeBERT shares the same tokenizer as CodeBERT and therefore exhibits the same distribution, so only the results for CodeBERT are shown. The figure indicates that the majority of samples contain fewer than 512 tokens for all smell types. Although some samples slightly exceed the 512-token boundary, almost all remain below 600 tokens. This suggests that truncation beyond 512 tokens affects only a small portion of the dataset.
Figure 10. Token length distribution for CodeBERT and CodeT5 across CC, CM, FE, and DC
To further evaluate whether the truncated segments contain information that is critical for code smell detection, we conducted an additional manual analysis using CodeBERT as a representative model. Before the main review, we performed a pilot annotation: for each code smell type, we randomly selected 10 positive and 10 negative samples, which were independently labeled by the first and fourth authors. Across all smell types, the two reviewers agreed on 85-95% of cases, indicating high consistency and supporting the reliability of our labeling. Following this pilot, we proceeded with the main manual review. For each code smell type, we randomly sampled 50 positive and 50 negative instances whose lengths exceed the 512-token limit. These samples were manually reviewed to determine whether the truncated portion included code elements that was essential for correct classification. Table 13 summarizes the results.
Table 13. Number of truncated samples in which the truncated portion does not affect the detection of CC, CM, FE, and DC
$C$ CM FE DC
#Positives #Negatives #Positives #Negatives #Positives #Negatives #Positives #Negatives
49 49 50 47 50 49 50 48
The results show that for nearly all sampled instances, truncation beyond 512 tokens does not remove information critical for code smell detection. Only a few cases (1–3 samples per category) potentially lose relevant information. Most essential indicators appear within the first 512 tokens, while the truncated segments typically contain auxiliary logic or repeated structures that do not affect classification.
Overall, these findings, combined with the token length statistics in Figure 10, suggest that context window limitations of CodeBERT, GraphCodeBERT, and CodeT5 have minimal impact on evaluation results. Nonetheless, we acknowledge that this review was conducted on a limited subset; for exceptionally long or complex methods, truncation could still obscure semantically important content. Future work could extend this analysis to a larger portion of the dataset to further quantify any residual effects more comprehensively.